OpenStudy (anonymous):

x(squared)+2x-3=0 solve by completing the square???

OpenStudy (anonymous):

Move the loose number over to the other side. \[x ^{2} +2x = 3\] Take half of the coefficient (don't forget the sign!) of the x-term, and square it. Add this square to both sides of the equation \[x ^{2} + 2x + 1 ^{2} = 3 +1^{2}\] Convert the left-hand side to squared form, and simplify the right-hand side. (This is where you use that sign that you kept track of earlier. You plug it into the middle of the parenthetical part.) \[(x+1)^{2} = 4\] Square-root both sides, remembering the"±" on the right-hand side. Simplify as necessary \[\sqrt{(x+1)^{2}} = \pm \sqrt{4}\] \[x+1 = \pm 2\] \[x = -1\pm2\] x = -3 or 1